A stocks returns have the following distribution: Demand for the Probability of this Rate of Return If Companys Products -
Answer #1

Expected Return = 0.1* - 0.28+ 0.1 * -0.13 + 0.4* 0.11 + 0.3* 0.28 + 0.1*0.63 = 0.15

SD of sample = Under root (X-Mean of X) / n-1

= 0.14

SD of population= Under root (X-Mean of X) / n

=0.12

Coefficient of Vartion = SD/Mean * 100

0.12/0.15 * 100 = 80

Sharpe ration = Return - Rf / SD = 0.15 - 0.04 / 0.12 = 0.91